OpenStudy (anonymous):

what does d= in this equation? d over five = 4

OpenStudy (anonymous):

d/5 = 4 d = 4*5 d = ?

OpenStudy (the_fizicx99):

20/5 = 4

OpenStudy (the_fizicx99):

>.> Ma bad, sorry.

OpenStudy (anonymous):

its fine and yeah thanks was looking at that thanks everyone